# Diagonalization $of$ Polynomial-Time Deterministic Turing Machines via Nondeterministic Turing Machines

### v31

@@ 3 Enumeration of All Polynomial-Time Deterministic Turing Machines
$\pi(k_{1},k_{2}):=\frac{1}{2}(k_{1}+k_{2})(k_{1}+k_{2}+1)+k_{2}$
where $k_{1},k_{2}\in\mathbb{N}$. Since the Cantor pairing function (see Figure 2 below which is from [A2]) is invertible (see [A2]), it is a bijection between $\mathbb{N}\times\mathbb{N}$ and $\mathbb{N}$. As we have shown that any polynomial-time deterministic Turing machine itself is an integer, we can place any polynomial-time deterministic Turing machine $M$ and its order $k$ to the tuple $(M,k)$ and use the Cantor pairing function to map the tuple $(M,k)$ to an integer in $\mathbb{N}_{1}$. Recall that there is a bijection between $\mathbb{N}$ and $\mathbb{N}_{1}$. Obviously, by Definition 5, the inverse of such a Cantor pairing function is an enumeration of the set $\{(M,k)\}$ 161616 Thus, the set $\mathbb{N}_{1}\times\mathbb{N}_{1}$ is enumerable or countable. of all polynomial-time deterministic Turing machines.
Figure 2: Cantor pairing function
4 Diagonalization against All Polynomial-Time Deterministic Turing Machines
We can now design a four-tape nondeterministic Turing machine $M_{0}$ which treats its input string $x$ both as an encoding of a tuple $(M,k)$ and also as the input to the polynomial-time deterministic Turing machine $M$. One of the capabilities possessed by $M_{0}$ is the ability to simulate a deterministic Turing machine, given its specification. By Lemma 2.2, the simulation can be done within
@@ 4 Diagonalization against All Polynomial-Time Deterministic Turing Machines
steps, so we shall have $M_{0}$ determine whether the deterministic $n^{k}+k$ time-bounded Turing machine $(M,k)$ accepts the input $x$ without using more than $O(T(n)\log T(n))$ steps where
@@
Proof. Let $M_{0}$ be a four-tape nondeterministic Turing machine which operates as follows on an input string $x$ of length of $n$.
  1. 1.
    $M_{0}$ decodes the tuple encoded by $x$. If $x$ is not the encoding of some polynomial-time deterministic Turing machine $\widehat{M}_{j}$ for some $j$ then GOTO $5$, else determines $t$, the number of tape symbols used by $\widehat{M}_{j}$; $s$, its number of states; and $k$, its order. The third tape of $M_{0}$ can be used as “scratch” memory to calculate $t$.
  2. 2.
@@
    cells each, the blocks being separated by single cell holding a marker ##\##, i.e., there are
@@
    $n^{k+1}.$
  4. 4.
@@
  5. 5.
    Since $x$ is not encoding of some deterministic Turing machine. Then $M_{0}$ sets up a block of
@@
    By using its nondeterministic choices, $M_{0}$ moves as per the path given by $x$. The moves of $M_{0}$ are counted in binary in the block of tape $3$. If the counter on tape $3$ overflows, then $M_{0}$ halts. $M_{0}$ accepts $x$ if and only if there is a computation path from the start state of $M_{0}$ leading to the accept state and the total number of moves can not exceed
    $n^{2},$
    steps, so is within $O(n)$. Note that the number of $2$ in
@@
Suppose now $L_{d}$ were accepted by some, say, the $i$-th deterministic Turing machine in the enumeration $e$ which is a deterministic
@@
time-bounded Turing machine $\widehat{M}_{i}$. Then by Corollary 11, we may assume that $\widehat{M}_{i}$ is a single-tape polynomial-time deterministic Turing machine. Let $\widehat{M}_{i}$ have $s$ states and $t$ tape symbols. Since $\widehat{M_{i}}$ appears infinitely often191919 We know that we may prefix $1$’s at will to find larger and larger integers representing the same set of quintuples of the same deterministic Turing machine $M_{i}$, thus there are infinitely binary strings of sufficiently long which represents deterministic Turing machine $M_{i}$. in the enumeration $e$, $M_{0}$ was to set the counter of tape $3$ to count up to $n^{k+1}$, and
@@
So, there exists a $N_{0}>0$ such that for any $N\geq N_{0}$,
$T(N)\log T(N)<N^{k+1}$
@@
Thus, on input $w$, $M_{0}$ has sufficient time to simulate $M_{w}$ and accepts if and only if $M_{w}$ rejects.202020 In the simulation of a polynomial-time deterministic Turing machine, $M_{0}$ only turns itself off mandatorily when the counter on tape $3$ overflows, i.e., its counter $\geq|w|^{k+1}.$ where $w$ is the input. But we assumed that $\widehat{M}_{i}$ accepted $L_{d}$, i.e., $\widehat{M}_{i}$ agreed with $M_{0}$ on all inputs. We thus conclude that $\widehat{M}_{i}$ does not exist, from which it immediately follows
@@
We need to remark more about item (18) in the proof of Theorem 9. Because the universal nondeterministic Turing machine $M_{0}$ diagonalizes against all polynomial-time deterministic Turing machines, it can flip the answer immediately when the simulation ends due to
@@
In fact, we can design our universal nondeterministic Turing machine $M_{0}$ more complicated. For example, since we can encode any polynomial-time nondeterministic Turing machine into a binary string representing itself followed by its order $110^{k}$ 1 (see [Lin21b]), the input to $M_{0}$ can be classified into three types: if the input is a polynomial-time deterministic Turing machine, then $M_{0}$ does the work specified in the proof of Theorem 9 except item (5); if the input is a polynomial-time nondeterministic Turing machine $N$ (of time complexity $n^{k}+k$), then $M_{0}$ can set the counter of tape $3$ to count up to
$n^{k+1}$
then to simulate $N$ and accept the input if and only if $N$ accepts; otherwise, $M_{0}$ rejects the input. Note that such design does not change $M_{0}$’s time complexity.
In general, the diagonalization technique used in the proof of Theorem 9 can not directly apply to diagonalizing against a collection of nondeterministic
@@
as observed by Cook [Coo73]. Thus, in his work [Coo73], Cook uses different techniques and then by “indirect” diagonalization, which is somewhat complicated (e.g. the ingredient of technique from [Iba72] with other complicated techniques), to show a nondeterministic time hierarchy theorem [Coo73] (it has to say that there exists other techniques such as lazy diagonalization that is also capable of showing nondeterministic time hierarchy theorem; see e.g. [AB09, For00, FS07, Zak83]). But fortunately, we diagonalize against all polynomial-time deterministic Turing machines rather than polynomial-time nondeterministic Turing machines in the Theorem 9 and its proof, so our nondeterministic Turing machine $M_{0}$ can flip the answer immediately when the simulation ends.
@@
Theorem 10
The universal nondeterministic Turing machine $M_{0}$ constructed in proof of Theorem 9 runs within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$.
Proof. The simplest way to show the theorem is to prove that for any input $w$ to $M_{0}$, there is a corresponding positive integer $i_{w}\in\mathbb{N}_{1}$ such that $M_{0}$ runs at most
@@
steps by the construction, so the corresponding integer $i_{x}$ is $1$ in this case (i.e., $i_{x}=1$). In both cases we have show that for any input $w$ to $M_{0}$, there is a corresponding positive integer $i_{w}\in\mathbb{N}_{1}$ such that $M_{0}$ runs at most $|w|^{i_{w}+1}$ steps. So $M_{0}$ is a nondeterministic
@@
time-bounded Turing machine for any $k\in\mathbb{N}_{1}$. By Lemma 10, there is a single-tape nondeterministic Turing machine $M^{\prime}$ equivalent to $M_{0}$ and $M^{\prime}$ runs within time
@@
5 Proof of $L_{d}\in\mathcal{NP}$
At the moment, since the exponent $k$ in $O(n^{k})$, which is the running time of $M_{0}$, can be any integer in $\mathbb{N}_{1}$, it is not so transparent for some readers that the language $L_{d}$ accepted by $M_{0}$ running in $O(n^{k})$ for all $k\in\mathbb{N}_{1}$ is in $\mathcal{NP}$. For example, see [For21], in which Fortnow claims that $M_{0}$ to run in polynomial-time it must run in time $O(n^{c})$ for a fixed $c$. But to diagonalize all polynomial-time deterministic Turing machines then $M_{0}$ needs time $O(n^{k})$ for all $k\in\mathbb{N}_{1}$, including $k>c$. He hence asserts that the author made a common mistake and this mistake is not fixable due to the result of [BGS75]. In view of the above, at this moment in time the most notable difficulty is that there is no fixed constant $c\in\mathbb{N}_{1}$ such that $M_{0}$ to run within time $O(n^{c})$. We had to overcome such obstacles to show the fact of the theorem below, and admittedly, overcoming such obstacles requires sophisticated observation from novel perspectives (however, there is a simple and fast way to get the intuition that $L_{d}\in\mathcal{NP}$, see footnote $26$, although the footnote $26$ is not considered a rigorous mathematical proof):
Theorem 11
212121 To be more accurate, we should discuss this theorem with a single-tape nondeterministic time-bounded Turing machine $\widehat{M}_{0}$ which is of time complexity $O(n^{2k})$ for any $k\in\mathbb{N}_{1}$ accepting the language $L_{d}$. But the final outcome is the same as $M_{0}$.
The language $L_{d}$ is in $\mathcal{NP}$ where $L_{d}$ is accepted by $M_{0}$ which runs within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$.
Proof. Let us first define the family of languages
@@ 5 Proof of $L_{d}\in\mathcal{NP}$
$L_{d}^{i}\overset{{\rm def}}{=}$ language accepted by $M_{0}$ running within time $O(n^{i})$ for fixed $i\in\mathbb{N}_{1}$.
@@
during the computation exceed $n^{i+1}$ steps,
which technically can be done by adding a new tape to $M_{0}$ as a counter to count up to
@@
Then by the construction of $M_{0}$, namely, for an arbitrary input $w$ to $M_{0}$, there is a corresponding integer $i_{w}\in\mathbb{N}_{1}$ such that $M_{0}$ runs at most
@@
steps (in other words, $M_{0}$ runs at most $n^{i+1}$ steps for any $i\in\mathbb{N}_{1}$ where $n$ the length of input, see Theorem 10 above), we have
@@
$L_{d}^{i}\subseteq L_{d}^{i+1},\quad\text{for each fixed $i\in\mathbb{N}_{1}$}$
@@
$w\in L_{d}^{i+1}.$
This gives that for any fixed $i\in\mathbb{N}_{1}$,
(2) $\begin{split}L_{d}^{1}\subseteq L_{d}^{2}\subseteq\cdots\subseteq L_{d}^{i}\subseteq L_{d}^{i+1}\subseteq\cdots\end{split}$
Note further that for any fixed $i\in\mathbb{N}_{1}$, $L_{d}^{i}$ is the language accepted by the nondeterministic Turing machine $M_{0}$ within time $O(n^{i})$, i.e., at most $n^{i+1}$ steps,232323 When $M_{0}$ turns itself off mandatorily when its moves made by itself during the computation exceed $n^{i+1}$ steps, it is a polynomial-time nondeterministic Turing machine, i.e., it is a nondeterministic $n^{i+1}+(i+1)$ time-bounded Turing machine. we thus obtain
(3) $\begin{split}L_{d}^{i}\in\text{NTIME}[n^{i}]\subseteq\mathcal{NP},\quad\text{for any fixed $i\in\mathbb{N}_{1}$}.\end{split}$
Now, (1) together with (2) and (3) easily implies
$L_{d}\in\mathcal{NP},$
as required.
5.1 The $2$nd Simpler Proof
In fact, we can prove Theorem 21 by contradiction, which is much more simpler. To do so, after obtaining the relations (1) and (2), we can assume that

### v33

@@ 3. Enumeration of All Polynomial-Time Deterministic Turing Machines
$\pi(k_{1},k_{2}):=\frac{1}{2}(k_{1}+k_{2})(k_{1}+k_{2}+1)+k_{2},$
where $k_{1},k_{2}\in\mathbb{N}$. Since the Cantor pairing function (see Figure 2 below, which is from [A2]) is invertible (see [A2]), it is a bijection between $\mathbb{N}\times\mathbb{N}$ and $\mathbb{N}$. As we have shown that any polynomial-time deterministic Turing machine itself is an integer, we can place any polynomial-time deterministic Turing machine $M$ and its order $k$ in the tuple $(M,k)$ and use the Cantor pairing function to map the tuple $(M,k)$ to an integer in $\mathbb{N}_{1}$. Recall that there is a bijection between $\mathbb{N}$ and $\mathbb{N}_{1}$. Obviously, by Definition 5, the inverse of such a Cantor pairing function is an enumeration of the set $\{(M,k)\}$161616 Thus, the set $\mathbb{N}_{1}\times\mathbb{N}_{1}$ is enumerable or countable. of all polynomial-time deterministic Turing machines.
Figure 2. Cantor pairing function
We can now design a four-tape nondeterministic Turing machine $M_{0}$ that treats its input string $x$ both as an encoding of a tuple $(M,k)$ and also as the input to the polynomial-time deterministic Turing machine $M$. One of the capabilities possessed by $M_{0}$ is the ability to simulate a deterministic Turing machine, given its specification. By Lemma 2.5, the simulation can be done within
@@ 4. Diagonalization against All Polynomial-Time Deterministic Turing Machines
steps, so we shall have $M_{0}$ determine whether the deterministic $n^{k}+k$ time-bounded Turing machine $(M,k)$ accepts the input $x$ without using more than $O(T(n)\log T(n))$ steps, where
@@
Proof 4.2.
Let $M_{0}$ be a four-tape nondeterministic Turing machine that operates as follows on an input string $x$ of length $n$.
  1. (1)
    $M_{0}$ decodes the tuple encoded by $x$. If $x$ is not the encoding of some polynomial-time deterministic Turing machine $\widehat{M}_{j}$ for some $j$, then GOTO $5$; else determine $t$, the number of tape symbols used by $\widehat{M}_{j}$; $s$, its number of states; and $k$, its order. The third tape of $M_{0}$ can be used as “scratch” memory to calculate $t$.
  2. (2)
@@
    cells each, the blocks being separated by a single cell holding a marker ##\##, i.e., there are
@@
    n^k+1.171717 Assume that $\widehat{M}_{j}$ is a deterministic $T(n)$ time-bounded Turing machine where $T(n)=n^{k}+k,$ then by Lemma 2.5, the simulation by $M_{0}$ can be done within time $T(n)\log T(n)$, which is less than $n^{k+1}$, so we set the counter to count up to $n^{k+1}$.
  4. (4)
@@
  5. (5)
    Since $x$ is not an encoding of some deterministic Turing machine. Then $M_{0}$ sets up a block of
@@
    By using its nondeterministic choices, $M_{0}$ moves as per the path described by $x$. The moves of $M_{0}$ are counted in binary in the block of tape $3$. If the counter on tape $3$ overflows, then $M_{0}$ halts. $M_{0}$ accepts $x$ if and only if there is a computation path from the start state of $M_{0}$ leading to the accept state and the total number of moves can not exceed
    $n^{2}$
    steps, so it is within $O(n)$. Note that the number of $2$ in
@@
Suppose now $L_{d}$ were accepted by some, say, the $i$-th deterministic Turing machine in the enumeration $e$, which is a deterministic
@@
time-bounded Turing machine $\widehat{M}_{i}$. Then by Corollary 11, we may assume that $\widehat{M}_{i}$ is a single-tape polynomial-time deterministic Turing machine. Let $\widehat{M}_{i}$ have $s$ states and $t$ tape symbols. Since $\widehat{M_{i}}$ appears infinitely often191919 We know that we may prefix $1$s at will to find larger and larger integers representing the same set of quintuples of the same deterministic Turing machine $M_{i}$; thus, there are infinitely many binary strings of sufficient length that represent deterministic Turing machine $M_{i}$. in the enumeration $e$, $M_{0}$ was to set the counter of tape $3$ to count up to $n^{k+1}$, and
@@
So, there exists an $N_{0}>0$ such that for any $N\geq N_{0}$,
$T(N)\log T(N)<N^{k+1},$
@@
Thus, on input $w$, $M_{0}$ has sufficient time to simulate $M_{w}$ and accepts if and only if $M_{w}$ rejects.202020 In the simulation of a polynomial-time deterministic Turing machine, $M_{0}$ only turns itself off mandatorily when the counter on tape $3$ overflows, i.e., its counter $\geq|w|^{k+1},$ where $w$ is the input. But we assumed that $\widehat{M}_{i}$ accepted $L_{d}$, i.e., $\widehat{M}_{i}$ agreed with $M_{0}$ on all inputs. We thus conclude that $\widehat{M}_{i}$ does not exist, from which it immediately follows
@@
We need to remark more about item (18) in the proof of Theorem 4.1. Because the universal nondeterministic Turing machine $M_{0}$ diagonalizes against all polynomial-time deterministic Turing machines, it can flip the answer immediately when the simulation ends due to
@@
In fact, we can design our universal nondeterministic Turing machine $M_{0}$ to be more complicated. For example, since we can encode any polynomial-time nondeterministic Turing machine into a binary string representing itself followed by its order $110^{k}$ 1 (see [Lin21b]), the input to $M_{0}$ can be classified into three types: If the input is a polynomial-time deterministic Turing machine, then $M_{0}$ does the work specified in the proof of Theorem 4.1 except item (5); if the input is a polynomial-time nondeterministic Turing machine $N$ (of time complexity $n^{k}+k$), then $M_{0}$ can set the counter of tape $3$ to count up to
$n^{k+1},$
then to simulate $N$ and accept the input if and only if $N$ accepts; otherwise, $M_{0}$ rejects the input. Note that such a design does not change $M_{0}$’s time complexity.
In general, the diagonalization techniques used in the proof of Theorem 4.1 can not directly apply to diagonalizing against a collection of nondeterministic
@@
as observed by Cook [Coo73]. Thus, in his work [Coo73], Cook uses different techniques and then, by “indirect” diagonalization, which is somewhat complicated (e.g., the ingredient of technique from [Iba72] with other complicated techniques), to show a nondeterministic time hierarchy theorem [Coo73] (it has to say that there exist other techniques such as lazy diagonalization that are also capable of showing nondeterministic time hierarchy theorems; see e.g., [AB09, For00, FS07, Zak83]). But fortunately, we diagonalize against all polynomial-time deterministic Turing machines rather than polynomial-time nondeterministic Turing machines in Theorem 4.1 and its proof, so our nondeterministic Turing machine $M_{0}$ can flip the answer immediately when the simulation ends.
@@
Theorem 4.4.
The universal nondeterministic Turing machine $M_{0}$ constructed in proof of Theorem 4.1 runs within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$.
Proof 4.5.
The simplest way to show the theorem is to prove that for any input $w$ to $M_{0}$, there is a corresponding positive integer $i_{w}\in\mathbb{N}_{1}$ such that $M_{0}$ runs at most
@@
steps by the construction, so the corresponding integer $i_{x}$ is $1$ in this case (i.e., $i_{x}=1$). In both cases we have shown that for any input $w$ to $M_{0}$, there is a corresponding positive integer $i_{w}\in\mathbb{N}_{1}$ such that $M_{0}$ runs at most $|w|^{i_{w}+1}$ steps. So $M_{0}$ is a nondeterministic
@@
time-bounded Turing machine for any $k\in\mathbb{N}_{1}$. By Lemma 10, there is a single-tape nondeterministic Turing machine $M^{\prime}$ equivalent to $M_{0}$, and $M^{\prime}$ runs within time
@@
5. Proof of $L_{d}\in\mathcal{NP}$
At the moment, since the exponent $k$ in $O(n^{k})$, which is the running time of $M_{0}$, can be any integer in $\mathbb{N}_{1}$, it is not so transparent for some readers that the language $L_{d}$ accepted by $M_{0}$ running in $O(n^{k})$ for all $k\in\mathbb{N}_{1}$ is in $\mathcal{NP}$. For example, see [For21], in which complexity theorist Fortnow claims that for $M_{0}$ to run in polynomial time, it must run in time $O(n^{c})$ for a fixed $c$. But to diagonalize all polynomial-time deterministic Turing machines, then $M_{0}$ needs time $O(n^{k})$ for all $k\in\mathbb{N}_{1}$, including $k>c$. He hence asserts that the author made a common mistake, and this mistake is not fixable due to the result of [BGS75]. In view of the above, at this moment in time the most notable difficulty is that there is no fixed constant $c\in\mathbb{N}_{1}$ such that $M_{0}$ runs within time $O(n^{c})$. We had to overcome such obstacles to show the fact of the theorem below, and admittedly, overcoming such obstacles requires sophisticated observation from novel perspectives (however, there is a simple and fast way to get the intuition that $L_{d}\in\mathcal{NP}$, see footnote $26$, although the footnote $26$ is not considered a rigorous mathematical proof):
Theorem 5.1.
212121 To be more accurate, we should discuss this theorem with a single-tape nondeterministic time-bounded Turing machine $\widehat{M}_{0}$ that is of time complexity $O(n^{2k})$ for any $k\in\mathbb{N}_{1}$ accepting the language $L_{d}$. But the final outcome is the same as $M_{0}$.
The language $L_{d}$ is in $\mathcal{NP}$, where $L_{d}$ is accepted by $M_{0}$, which runs within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$.
Proof 5.2.
Let us first define the family of languages
@@ 5. Proof of $L_{d}\in\mathcal{NP}$
$L_{d}^{i}\overset{{\rm def}}{=}$ language accepted by $M_{0}$ running within time $O(n^{i})$ for a fixed $i\in\mathbb{N}_{1}$.
@@
during the computation exceed $n^{i+1}$ steps.
Note that the above definition of language $L_{d}^{i}$ technically can be done by adding a new tape to $M_{0}$ as a counter to count up to
@@
Then by the construction of $M_{0}$, namely, for an arbitrary input $w$ to $M_{0}$, there is a corresponding integer
$i_{w}\in\mathbb{N}_{1}$
such that $M_{0}$ runs at most
@@
steps (in other words, $M_{0}$ runs at most $n^{i+1}$ steps for any $i\in\mathbb{N}_{1}$ where $n$ is the length of the input; see Theorem 4.4 above), we have
@@
$L_{d}^{i}\subseteq L_{d}^{i+1},\quad\text{for each fixed $i\in\mathbb{N}_{1}$,}$
@@
5.1. The $2$nd Simpler Proof
In fact, we can prove Theorem 21 by contradiction, which is much simpler. To do so, after obtaining the relations (1) and (LABEL:eq2), we can assume that